OpenStudy (anonymous):

Express the function f(x)=sqrt(18x^2+8) as a composition f(x)=f(g(x)) of two simpler functions f(x) and g(x)

OpenStudy (anonymous):

Usually you would make g(x) the function inside the square root sign.

OpenStudy (anonymous):

I still do not understand. g(x)=18x^2+8 ? That's not the right answer

OpenStudy (anonymous):

Are you sure because.\[h(x) = f(g(x)) \\g(x) = 18x^2 +8 \\f(x) = \sqrt(g(x)) \]